OpenStudy (anonymous):

match them up :) exponential decay recursive sequence geometric sequence exponential growth arithmetic sequence with 2, 9, 16, 23... 3, 15, 75, 375... 0, 4, 4, 8... y = (1. 1)x y = (1/4) x

OpenStudy (ranga):

What patten do you see in this sequence: 2, 9, 16, 23...? Hint: Take the difference of successive numbers. That is, 9 -2 = ? 16 - 9 = ?, 23 - 16 = ?

OpenStudy (anonymous):

I know the difference in them but don't know the different names they match up with

OpenStudy (ranga):

In the first sequence the difference between two successive term is always the same and is equal to 7. This is an arithmetic sequence.

OpenStudy (ranga):

Take the second sequence: 3, 15, 75, 375... What is the pattern that you see? How can you get the second term from the first term? How can you get the third term from the second term?

OpenStudy (anonymous):

\[y = (1. 1)^x\]is exponential on account of the \(x\) in the exponent since \(1.1>1\) this is exponential growth